OpenStudy (anonymous):

What is the eccentricity of the graph of (x-2)^2/121 - (y+6)^2/64=1?

OpenStudy (loser66):

wow, it's cal2 I need time to look up at my note, friend.

OpenStudy (anonymous):

thats alright :)

OpenStudy (ness9630):

Is this really calc 2?

OpenStudy (ness9630):

I thought this was like Alg II

OpenStudy (anonymous):

Well it's suppose to be alg 2 lol

OpenStudy (ness9630):

@Loser66 I think you're getting a little too ahead of yourself friend xD

OpenStudy (loser66):

@DebbieG here, Let her do it, Please @DebbieG

OpenStudy (ness9630):

This is eccentricity of an ellipse, right?

OpenStudy (anonymous):

I believe so

OpenStudy (debbieg):

lol @Loser66 thanks for your faith in me but I'm not familiar with this topic! Was just looking it up.... http://www.mathsisfun.com/geometry/eccentricity.html

OpenStudy (ness9630):

Isn't it this:

OpenStudy (loser66):

Ness and the asker said that it's alg and I never took alg before. I just mess it up with my cal2 hehehe.. so, I don't know how to solve @DebbieG

OpenStudy (ness9630):

Come on loser, you know me bud ;P

OpenStudy (ness9630):

Anyways, i gtg, tell me how it goes!

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (debbieg):

http://www.mathsisfun.com/geometry/hyperbola.html If I'm interpreting this correctly, then isn't it: \[\Large e=\frac{ \sqrt{121+64} }{11 }=\frac{ \sqrt{185} }{11 }\]

OpenStudy (loser66):

hahaha, it's exactly the answer from my mess but they said that it's not cal2

OpenStudy (anonymous):

its algebra 2 lol and thank you for helping, you guys are great :)

OpenStudy (loser66):

@DebbieG they said that it's an ellipse. I passed out

OpenStudy (debbieg):

LOL well, like I said, this isn't a topic I'm very familiar with but it seems like pretty straight forward application of the formula. :) It's a hyperbola.... Wolfram said so! ;)

OpenStudy (loser66):

ooooh wolfram said so? hehehe... a little bit confidence is back to me. I solved it by looking up my note, not from that tool. OOOOOK , no matter what happen. I .... give up.